OpenStudy (anonymous):

if f and g are differentiable at x then so is their product. (f.g)'(x)=f(x)g'(x)+g(x)f'(x) lgbasallote plz the answer

OpenStudy (anonymous):

this is a fact, not a question

OpenStudy (lgbasallote):

maybe the question is to prove?

OpenStudy (anonymous):

depends on what you know how easy this is. it is pretty straightforward to prove from the definition that \[(f^2)'=2ff'\] then you can write the product as squares and do some minor algebra to get it

OpenStudy (anonymous):

yes the prove igbasallote

OpenStudy (anonymous):

what ths prove

OpenStudy (anonymous):

i need prove lgbasallote

OpenStudy (anonymous):

you can start by knowing that \[(f^2)'=2ff'\] which you can prove by the definition easily. then note that \[(f+g)^2=f^2+2fg+g^2\] so \[fg=\frac{1}{2}((f+g)^2-f^2-g^2)\]

OpenStudy (anonymous):

take the derivative of the right hand side using the rule for squares to find the derivative of the left hand side

OpenStudy (anonymous):

lgbasallote where r u ?

OpenStudy (lgbasallote):

the proof of this thing is very confusing...but that is as simple as it can be ^

OpenStudy (anonymous):

ok

OpenStudy (lgbasallote):

expanding what satellite73 said... fg = 1/2(2(f+g)(df + dg) - 2fdf -2gdg) you can cancel the 1/2 there... so... fg = (f+g)(df+dg) - (df + dg) fg =fdf +gdf +fdg +gdg - df - dg fdf,gdg,df and dg are all equal to 1 if i'm not mistaken...so... fg = 1 + gdf + fdg + 1 -1 -1 cancel... fg = gdf + fdg thus... the product rule :D

OpenStudy (zarkon):

these things are on wikipedia http://en.wikipedia.org/wiki/Product_rule#Proof_of_the_product_rule